8. Max’s Christmas tree lot has a monopoly on sales of Christmas trees. To increase his sales from 100 trees to 108 trees, he must drop the price of all his trees from R28 to R26. What is Max’s marginal revenue when he lowers his price and increases his sales from 100 to 108 trees?  [1] R2,808 [2] R28 [3] R26 [4] R8 [5] R1 


1
Expert's answer
2021-04-13T07:16:40-0400
"\frac {108\times 26 -100\times 28}{108-100}=1"

Answer [5]